Divisional Income Statements and Return on Investment Analysis

E.F. Lynch Company is a diversified investment company with three operating divisions organized as investment centers. Condensed data taken from the records of the three divisions for the year ended June 30, 20Y8, are as follows:


Mutual Fund Division

Electronic Brokerage Division

Investment Banking Division
Fee revenue $1,670,000 $1,740,000 $1,620,000
Operating expenses 901,400 730,200 1,224,000
Invested assets 6,100,000 5,100,000 3,300,000

The management of E.F. Lynch Company is evaluating each division as a basis for planning a future expansion of operations.

Required:

1. Prepare condensed divisional income statements for the three divisions, assuming that there were no service department charges.

E.F. Lynch Company
Divisional Income Statements
For the Year Ended June 30, 20Y8
Mutual Fund Division Electronic Brokerage Division Investment Banking Division
Fee revenue $ $ $
Operating expenses
Income from operations $ $ $

2. Using the DuPont formula for rate of return on investment, compute the profit margin, investment turnover, and rate of return on investment for each division. Round your answers to one decimal place.

Division Profit Margin Investment Turnover ROI
Mutual Fund Division % %
Electronic Brokerage Division % %
Investment Banking Division % %

3. When faced with limited funds for expansion, management should consider an expansion of the   Division first.
Answer #1

1.

E.F. Lynch Company
Divisional income statements
For the year ended June 30,20Y8
Mutual Fund division Electronic Brokerage Division Investment Banking Division
Fee revenue 1670000 1740000 1620000
Operating expenses -901400 -730200 -1224000
Income from operations $768600 1009800 396000

2.

Mutual Fund Division

Profit margin = Income from operations/Fee revenue

= 768600/1670000

= 46.02%

Investment turnover = Fee revenue/Invested assets

= 1670000/6100000

= 0.27

ROI = Income from operations/Invested assets

= 768600/6100000

= 12.6%

Electronic Brokerage Division

Profit margin = Income from operations/Fee revenue

= 1009800/1740000

= 58.03%

Investment turnover = Fee revenue/Invested assets

= 1740000/5100000

= 0.34

ROI = Income from operations/Invested assets

= 1009800/5100000

= 19.80%

Investment Banking Division

Profit margin = Income from operations/Fee revenue

= 396,000/1,620,000

= 24.44%

Investment turnover = Fee revenue/Invested assets

= 1,620,000/3,300,000

= 0.49

ROI = Income from operations/Invested assets

= 396,000/3,300,000

= 12%

Division Profit margin Investment turnover ROI
Mutual Fund Division 46.02% 0.27 12.60%
Electronic Brokerage Division 58.03% 0.34 19.80%
Investment Banking Division 24.44% 0.49 12%

3.

Since ROI of Electronic Brokerage Division is highest, hence in case of shortage of funds, it should be expanded first.